OpenStudy (anonymous):

Which of the following is a solution to the equation y = 3x – 1? (1 point) (4, 1) (2, 5) (4, 3) (0, –3)

OpenStudy (anonymous):

To be honest, you can easily figure this out by plugging in the x and y values in to the equation. If both sides are equal then it is a solution. y=3x-1 1=3(4)-1. 1=12-1 1=11 => (4,1) not a solution 5=3(2)-1 5=6-1 5=5 => (2, 5) is a solution 3=3(4)-1 3=12-1 3=11 => (4, 3) not a solution -3=3(0)-1 -3=-1 => (0, –3) not a solution

OpenStudy (anonymous):

oh I got it .